### v2

# Distillation Improves Visual Place Recognition for Low Quality Images

Abstract
Real-time visual localization often utilizes online computing, for which query images or videos are transmitted to remote servers for visual place recognition (VPR). However, limited network bandwidth necessitates image-quality reduction and thus the degradation of global image descriptors, reducing VPR accuracy. We address this issue at the descriptor extraction level with a knowledge-distillation methodology that learns feature representations from high-quality images to extract more discriminative descriptors from low-quality images. Our approach includes the Inter-channel Correlation Knowledge Distillation (ICKD) loss, Mean Squared Error (MSE) loss, and Triplet loss. We validate the proposed losses on multiple VPR methods and datasets subjected to JPEG compression, resolution reduction, and video quantization. We obtain significant improvements in VPR recall rates under all three tested modalities of lowered image quality. Furthermore, we fill a gap in VPR literature on video-based data and its influence on VPR performance. This work contributes to more reliable place recognition in resource-constrained environments. Source code and data available at https://github.com/ai4ce/LoQI-VPR.
I Introduction
Visual Place Recognition (VPR) has rapidly become a foundational component in machine perception. By condensing RGB images into compact vector representations, devices can recognize previously visited locations and recall their locations through only vision input and an image database of known locations [1, 2]. VPR has become a critical component in applications ranging from robots navigating complex industrial settings to augmented reality systems overlaying digital data onto the physical world to wayfinding applications that support vulnerable populations [3, 4, 5].
The adaptability of VPR to diverse environments has expanded its use cases. From smartphones to autonomous vehicles, the need for efficient and precise VPR systems is rising. However, this demand, especially in consumer devices, introduces novel challenges. A primary concern is the rapid growth in scene diversity and volume for real-world VPR applications, leading to large image databases that pose storage challenges for devices with limited capacity [6, 7].
Furthermore, the goal of more precise image descriptor vectors and thus better VPR performance has led to more complex VPR methods [8, 9, 10]. While such advancements do not delay descriptor matching, they increase the computational demands of extracting descriptors from images. Thus, real-time VPR on standard devices becomes more challenging due to significant computational cost [11] in addition to storage needs.
Figure 1: Enhanced Low Quality Image Retrieval: After knowledge distillation from high quality images, multiple VPR methods succeeds in their VPR recall of this JPEG-compressed query image from the Mapillary SLS dataset, while the same methods fail to do so beforehand.
Transitioning to online computation, where images are transmitted over the network for centralized processing, could be a solution for both issues but not without difficulties. While many devices simultaneously capture images to query the VPR image database, real-time communication with servers becomes a bottleneck. As most real-world VPR applications localizes over image sequences, the underlying video streaming service could alleviate network bandwidth pressure through reducing resolution and increasing compression of images and quantization of video streams. However, these measures degrade image descriptors and reduce localization accuracy [12].
For this online VPR paradigm, we propose a VPR-specific knowledge-distillation framework that improves descriptor generation and thus VPR performance under low-quality images, offering the following contributions:
  • We develop a strategy to distill state-of-the-art VPR methods’ knowledge of high-quality images into that of low-quality images, significantly improving performance on low-quality images.
  • We demonstrate three distillation loss functions on multiple public datasets and analyze the nature of each on descriptor quality.
  • We validate the generalizability of our approach by extending it to multiple forms of image quality reduction, including resolution decrease and video quantization, curating a video-based VPR dataset in the process to also fill a gap in video data for VPR research.
II Related Work
II-A VPR’s Challenges with Low-Quality Images
Prevailing VPR literature frames the problem of image-based localization as an image retrieval task [2]. Under this modality, VPR performance is heavily dependent on image descriptors’ abilities to relate images with distinct appearances to geographical proximity. Earlier VPR techniques such as GIST [13], HOG [14], and VLAD [15] generate compact image descriptors through hand-crafted local feature extractors. For better image matching performance, NetVLAD [8] and other recent VPR methods such as Patch-NetVLAD [16], MixVPR [9], AnyLoc [10], DINOv2 SALAD [17], Bag-of-Queries [18], and CricaVPR [19] leverage progress in deep learning and utilize convolutional and attention-based neural networks to produce learned local features and extract global image-level descriptors.
However, the rising complexity of VPR global descriptor extraction, while improving image retrieval accuracy, poses computational challenges, particularly in resource-constrained devices across real-world applications. Although special-purpose VPR techniques [11, 20, 21] with low computational cost address this limitation, the transmission of images for centralized descriptor extraction is often still necessary for deploying VPR techniques at scale [22, 23], completely free of both storage and computational constraints. Nevertheless, the resulting decrease in image quality from networking limitations reduces localization accuracy, which occurs under both lowered image resolution [12] and increased JPEG compression [24]. As many existing VPR datasets consist of purely JPEG images [25, 26, 27, 28, 29, 30, 31, 32, 33], the JPEG format’s prevalence both in VPR and in general [34] makes JPEG compression a main avenue of image quality reduction, besides decreasing image resolution. To our best knowledge, there exists a gap in VPR literature on publicly available datasets offering video data, and the effects of video quantization on VPR performance remain largely unexplored.
II-B Knowledge Distillation and Transfer Learning in VPR
Likewise, research on effective methods to extract more representative global descriptors from low-quality images is also lacking in the VPR community. Existing solutions have addressed the issue at the descriptor matching and image-retrieval level. The authors of [24] propose a sequence-based image-retrieval strategy [35], which queries the database images multiple times with lower per-query cost thanks to JPEG compression. Furthermore, predicting the retrieval success for each query image could favor points within the sequential queries that are more easily localized [36]. We instead wish to explore knowledge distillation as a means to directly improve descriptor extraction and thus increase the accuracy of database retrieval via individual queries.
After knowledge distillation was first proposed in [37], it has been applied to various computer-vision tasks such as super-resolution [38], image classification [39, 40], and segmentation [40]. More recently, knowledge distillation has also been used specifically in VPR to produce more discriminative global descriptors, such as learning from structural knowledge of an image-segmentation model [41]. We believe that knowledge distillation could similarly improve VPR accuracy for low-quality images.
Figure 2: Proposed Knowledge Distillation Methodology: Through ICKD loss between latent codes $z^{h}$ and $z^{l}$, MSE loss between global descriptors $v^{h}$ and $v^{l}$, and triplet loss on $v^{l}$, the student branch $f_{s}$ learns the teacher branch $f_{t}$’s knowledge of high quality images $I^{h}$ to extract more discriminative, representative $v^{l}$ from low quality images $I^{l}$.
III Distillation Methodology
Our goal was to enhance existing VPR methods to extract more representative global descriptors when given lower-quality images. Thus, we apply knowledge distillation techniques at the descriptor extraction level of VPR. For any VPR method, we aim to allow a student descriptor extractor handling low-quality images to approximate the output of its teacher counterpart, handling high-quality images.
Conceptual Model of Distillation for VPR
Given some VPR method, we initialize it twice to form the teacher branch $f_{t}$ and student branch $f_{s}$, processing high-quality image $I^{h}$ and its low-quality counterpart $I^{l}$ respectively. As illustrated by fig. 2, we conceptualize the teacher model as consisting of a feature encoder $f^{\theta}_{t}$ and descriptor aggregator $g^{\phi}_{t}$. Likewise, the student consists of a structurally identical encoder $f^{\theta}_{s}$ and aggregator $g^{\phi}_{s}$. Between the encoder and aggregator, the student and teacher branch generates 3D latent code $z^{l}$ and $z^{h}$, respectively. Finally, both branches aggregate their latent codes into global descriptors $v^{l}$ and $v^{h}$. Our distillation methodology has three loss functions that operate on this understanding of VPR methods.
ICKD Loss
The Inter-Channel Correlation Knowledge Distillation (ICKD) loss [40] operates on the latent codes $z^{l}$ and $z^{h}$. Given the potential size differences between inputs $I^{l}$ and $I^{h}$, the latent codes $z^{l}$ and $z^{h}$ may differ in dimensions. Traditional distillation losses, which typically require matching dimensions for the latent codes, are not directly applicable. Instead, the ICKD loss computes a standard-size Inter-Channel Correlation (ICC) matrix for both $f^{\theta}_{s}$ and $f^{\theta}_{t}$, allowing the former to approximate the latter.
To illustrate ICKD loss, define $I^{l}$ with resolution $W\times H$ and let the dimensions of their latent codes $z^{h}$ and $z^{l}$ be $c\times W^{\prime}\times H^{\prime}$ and $c\times w^{\prime}\times h^{\prime}$, respectively. Leveraging $c$ being fixed, the ICKD concatenates along the remaining dimensions $w^{\prime}$ and $h^{\prime}$ to produce vector $p^{l}$ with dimensions $c\times w^{\prime}\cdot h^{\prime}$. $p^{l}$ is then normalized along its second dimension, producing $\hat{p^{l}}=\frac{p^{l}}{\|p^{l}\|}$. The subsequent ICC matrix $C^{l}=\hat{p^{l}}\times\hat{p^{l}}^{T}$ with dimensions $c\times c$ is normalized to $\hat{C^{l}}=\frac{C^{l}}{\|C^{l}\|}$. A similar procedure for $z^{h}$ yields $\hat{C^{h}}\in\mathbb{R}^{c\times c}$. The ICKD loss, formulated to ensure $f^{\theta}_{s}$ learns from $f^{\theta}_{t}$, is defined as:
$L_{\theta 1}=\|\hat{C^{l}}-\hat{C^{h}}\|_{2}.$ (1)
Mean Squared Error (MSE) Loss
In addition to comparing latent codes, we directly compute the discrepancies between the image descriptors $v^{l}$ and $v^{h}$. The MSE loss is expressed as:
$L_{\theta 2}=\|v^{l}-v^{h}\|_{2}^{2}.$ (2)
The MSE loss trains $f_{s}$ to produce similar global descriptors as $f_{t}$ despite different input images, gaining insights from $f_{t}$ by directly emulating $f_{t}$’s outputs [42].
Weakly Supervised Triplet Ranking Loss
Applied exclusively within the student branch $f_{s}$, this loss is inspired by NetVLAD [8]. It ensures that the input image’s descriptor $v^{l}$ is more similar to that of a geographically proximate (positive) image, $\{v^{l}_{p}\}$, than that of a distant (negative) image, $\{v^{l}_{n}\}$. The goal is to satisfy the inequality $d_{\theta}(v^{l},v^{l}_{p})<d_{\theta}(v^{l},v^{l}_{n})$, where $d_{\theta}$ denotes the Euclidean distance between descriptors.
The triplet loss is articulated as:
$L_{\theta 3}=\sum_{j}l\left(\min_{i}d_{\theta}^{2}(v^{l},{v^{l}_{p}}_{i})-d_{\theta}^{2}(v^{l},{v^{l}_{n}}_{j})+m\right).$ (3)
In this equation, $l$ represents the hinge loss, defined as $l(x)=\max(x,0)$. The term $m$ is a margin parameter, ensuring a buffer between positive and negative pairs, thereby enhancing the discriminative capabilities of the learned embeddings. The component $\min_{i}d_{\theta}^{2}(v^{l},{v^{l}_{p}}_{i})$ selects the descriptor of the most similar image as estimated by the student model $f_{s}$, ensuring the loss is attuned to the nearest matches.
Composite Loss
We also aim to assess the effectiveness of a combined loss function relative to the individual application of the aforementioned losses. We compute a cumulative loss as:
$L=L_{\theta_{1}}+\alpha L_{\theta_{2}}+\beta L_{\theta_{3}}.$ (4)
In this equation, $\alpha$ and $\beta$ are weighting coefficients for $L_{\theta 2}$ and $L_{\theta 3}$, balancing the influence of each loss within the composite function. Our experiments will also examine each loss’s individual effects on distillation.
IV Experimental Setup
To analyze the efficacy of $L_{\theta 1}$, $L_{\theta 2}$, $L_{\theta 3}$, and $L$, we performed distillation on the student branch $f_{s}$ using the GSV-Cities VPR dataset [32]. Afterwards, we evaluated $f_{s}$’s retrieval accuracy via the Deep Visual Geo-localization Benchmark [43] and some of its available testing datasets. For the following training and testing datasets, we define low quality images $I^{l}$ as their original images ($I^{h}$) after $90$% JPEG compression (quality level = $10$).
Following Tomita et al. [35], we evaluate $f_{s}$ with $I^{l}$ VPR database and queries to avoid arbitrarily hampering descriptor matching from inconsistent compression levels, which also confirms the advantage of decreased database building cost when expecting $I^{l}$ queries. Here, we reduce image quality via JPEG compression due to the absence of video data and other image formats. We would explore our methodology’s generalizability to low image resolution and high video quantization in section VI.
IV-A Datasets Configurations
IV-A1 Training Dataset
The GSV-Cities dataset contains approximately $530,000$ street view images of cities around the world collected from various years. The $23$ total cities in the dataset are divided into a higher and lower resolution group containing images at $640\times 480$ and $400\times 300$. To decrease training cost, we designated only the cities with lower resolution images to be used, effectively shrinking the dataset to approximately $194,000$ images. This selective use of GSV-Cities additionally investigates whether knowledge distillation could be successful without high quality training data in the absolute sense.
Table I: Image Count of Testing Datasets
Partition Mapillary SLS Nordland Tokyo 24/7
Database 18871 27592 75984
Queries 11084 2760 315
IV-A2 Testing Datasets
We tested $f_{s}$ after distillation using each loss combination by calculating its VPR recall rate on three VPR datasets organized using VPR Datasets Downloader, namely Mapillary SLS [29], Nordland [26], and Tokyo 24/7 [28]. For each dataset, we follow the designated test partition and its database and query image specifications. Additional context for each dataset is provided below:
  • Mapillary SLS is specifically curated for lifelong VPR featuring images from 30 cities.
  • Nordland is created with forward-facing recordings across four seasons from trains along a rail route. Our version of the dataset is sourced identically as [9].
  • Tokyo 24/7 depicts urban scenes at different times of day within Tokyo’s Shibuya ward.
Table II: Best Performance of VPR Methods after Distillation: For low quality images, the loss combination producing the highest recall rate is compared against the baseline of fine-tuning. For both distilled and fine-tuned weights, the change in VPR recall is represented as a delta relative to the performance of each method using pretrained weights. Within each dataset and method, greene text indicates the greatest improvement for every R@N, whereas any decrease relative to pretrained performance is marked as red. The recall rates using pretrained weights on unmodified $I^{h}$ images are provided as reference.
VPR Methods Configuration Mapillary SLS Nordland Tokyo 24/7
R@1 R@2 R@5 R@10 R@1 R@2 R@5 R@10 R@1 R@2 R@5 R@10
MixVPR pretrained ($I^{h}$) 82.73 86.67 89.73 91.65 57.79 64.13 71.49 76.41 87.30 89.52 92.06 93.65
pretrained ($I^{l}$) 71.87 76.61 81.22 84.19 31.05 36.12 44.13 51.23 66.03 73.33 78.73 82.22
finetuned +4.43 +3.86 +3.60 +2.99 +13.37 +14.28 +14.64 +14.46 +9.52 +7.62 +6.03 +5.40
ICKD +4.33 +3.69 +3.20 +2.93 +15.00 +16.59 +16.78 +15.94 +8.25 +6.03 +6.35 +6.67
CricaVPR pretrained ($I^{h}$) 74.74 80.92 86.09 88.75 87.64 90.65 94.24 95.80 90.16 92.38 95.56 96.19
pretrained ($I^{l}$) 68.14 74.41 80.24 83.19 63.51 69.71 77.46 82.64 74.29 79.68 84.76 87.94
finetuned -0.12 -0.64 -0.77 -0.23 -2.97 -2.93 -2.97 -2.79 +5.08 +4.44 +2.54 +1.90
ICKD +0.89 +1.39 +1.11 +1.14 +11.63 +10.80 +8.95 +7.03 +6.98 +5.08 +6.03 +4.76
DINOv2 SALAD pretrained ($I^{h}$) 89.20 92.40 94.70 95.84 88.08 90.94 94.13 95.98 97.14 97.46 98.73 99.05
pretrained ($I^{l}$) 84.60 88.85 91.89 93.68 67.90 73.88 80.40 84.24 89.21 92.06 95.87 96.51
finetuned -0.22 -0.57 -0.60 -0.76 +0.04 -0.40 -0.76 -0.04 -0.63 0.00 -0.95 0.00
ICKD +0.52 +0.45 +0.20 +0.30 +1.56 +1.81 +1.09 +1.20 +1.59 +1.27 +0.32 0.00
NetVLAD pretrained ($I^{h}$) 49.29 55.24 62.07 67.29 5.51 6.67 8.62 11.38 60.63 63.81 69.21 74.29
pretrained ($I^{l}$) 32.60 37.67 44.92 50.13 1.99 3.01 4.93 7.07 27.94 33.02 41.90 47.94
finetuned +0.03 +0.01 -0.02 -0.01 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
MSE +4.46 +5.32 +5.09 +5.01 +0.54 +0.43 -0.18 -0.72 +8.25 +6.35 +6.03 +5.08
AnyLoc pretrained ($I^{h}$) 56.51 62.31 68.28 72.89 12.90 16.23 20.18 24.28 88.25 91.11 94.92 96.83
pretrained ($I^{l}$) 48.04 56.45 63.84 69.17 10.11 12.97 17.86 21.92 83.49 87.30 91.75 95.87
ICKD + Triplet +0.99 +1.54 +2.68 +2.44 +1.74 +1.88 +1.52 +2.54 -5.71 -1.59 +1.59 -0.95
IV-B VPR Methods
To validate the applicability of knowledge distillation for VPR, we carried out the aforementioned experiments on five VPR methods from various points in the VPR literature, namely NetVLAD [8], MixVPR [9], AnyLoc [10], DINOv2SALAD [17], and CricaVPR [19]. While the former two methods have CNN backbones, the remaining methods are based on DINOv2 [44]. The diverse architectures of these five methods allow us to explore our distillation approach’s applicability to diverse VPR methodologies. Furthermore, to study each loss component’s effects on $f_{s}$’s performance, we trained $f_{s}$ under each of the seven possible combinations of $L_{\theta 1}$, $L_{\theta 2}$, and $L_{\theta 3}$, using at least one term from (4). For each VPR method, we acquire pretrained model weights released with their public source codes to act as $f_{t}$ and initialize $f_{s}$ for faster training convergence.
As a baseline to be compared against distillation, we also fine-tune all methods except for AnyLoc using their unmodified training source code on the same $I^{l}$ and pretrained weights as distillation. Given that AnyLoc uses off-the-shelf DINOv2 backbone and does not learn its aggregator weights, we only apply distillation to this method.
IV-C Experimental Setup
IV-C1 Training Configuration
We set the loss weighting coefficients in (4) to $\alpha=10^{5}$ and $\beta=10^{4}$. The learning rate is initialized at $10^{-5}$, modulated by a decay factor of $2\times 10^{-11}$, and further adjusted by an exponential rate of $0.99999$. Each training experiment (ie. one of seven loss combinations with one of five methods) consists of one complete iteration through the shrunken GSV-Cities dataset IV-A1 (ie. one epoch). For the triplet loss (3), we sampled $5$ negative images to form $v^{l}_{n}$ for each training image.
IV-C2 VPR Performance Evaluation
To assess $f_{s}$’s VPR performance after distillation, we use the metric Recall at N (referred to as R@N) found in [8, 9, 16, 7, 17, 10, 19] to calculate $f_{s}$’s recall rate, which measures the percentage of successfully localized query images. Given a choice of $N$, a query to the VPR database is deemed successful if one of the top $N$ retrieved database images is within $d$ meters of the query’s true position for some predetermined threshold $d$. The percentage of successfully retrieved query images becomes $f_{s}$’s R@N score for that $N$. In our evaluation procedures, we selected $N$ values of $1$, $2$, $5$, and $10$, whereas $d$ remained fixed at 25 meters following the distance thresholds used by the VPR methods we experimented with.
Figure 3: Recall Rates for All Losses: For each method, we compare $f_{s}$ after distillation with loss combinations as well as pretrained weights. Different marker shapes indicate different combinations of ICKD and MSE losses, whereas including the triplet loss is shown with a dashed line. For readability, the y-axis of each subplot is independently scaled.
V Results and Discussions
For each loss combination, We use $f_{s}$ after distillation to generate $v^{l}$ for the selected testing datasets IV-A2. We then calculate recall rates for different versions of $f_{s}$ and analyze the efficacy of each configuration of our proposed distillation methodology.
V-A Quantitative Results
Table II demonstrates the applicability of (4) and its components to a representative selection of VPR methods and datasets. Albeit with varying magnitudes, the deltas of recall rates under $v^{l}$ after distillation and those under $v^{l}$ from pretrained weights are generally positive. However, fine-tuning is inconsistent and mostly less effective than distillation. The Nordland recall rates for MixVPR and CricaVPR, two architecturally distinct methods, exemplifies distillation’s superior efficacy and applicability. There are only two instances where distillation partially failed to increase VPR recall rates under $v^{l}$ for some method and dataset, namely NetVLAD on Nordland and AnyLoc on Tokyo 24/7. As AnyLoc’s recall rates fluctuate greatly on Tokyo 24/7, we postulate the combination of the natures of this method and dataset to be an edge case. Since NetVLAD’s improvement on Nordland for all R@N is less than that on other datasets, we reserve this phenomenon for further analysis in V-B.
Besides numerically demonstrating the most effective distillation losses, we graphically compare each loss combination in fig. 3, where we analyze full recall rate data for all variants of $f_{s}$. Previously, we have observed that ICKD (1) produces best VPR performances for MixVPR and CricaVPR and that MSE (2) is best for NetVLAD, and these three methods benefit the most from distillation. Now, the relative superiority of the two aforementioned losses could be seen on all methods, where the recall rates for $f_{s}$ trained under either or both of them generally trend towards the top of all loss combinations. Here, AnyLoc and NetVLAD on Nordland are exceptions, where augmenting (1) or (2) with triplet loss (3) produces a better-performing $f_{s}$. However, besides NetVLAD, whose original training scheme uses the triplet loss, (3) is generally the least effective when used alone.
V-B Qualitative results
We further investigate distillation’s effects on MixVPR, CricaVPR, and NetVLAD, the most improved methods, and potentially account for the latter’s more limited improvement on Nordland. In fig. 4, we plot activation maps for each method’s encoder $f^{\theta}_{s}$ after distillation with their respective best-performing loss. From each of the three VPR testing datasets IV-A2, we visualize an $I^{l}$ query where $f_{s}$ improves upon $f_{t}$. On Mapillary SLS and Nordland, distillation mitigates the distraction of uninformative sky features especially for NetVLAD and MixVPR. More generally, distillation shifts focus from repetitive foreground features, such as road surfaces and ground-level vegetation, to more structurally distinct whole-scene features, such as the contours of buildings and farther tree lines. In the case of NetVLAD, the highly selective focus of $f^{\theta}_{s}$ could account for its lesser recall rate improvement on Nordland. For the other two methods, while MixVPR’s improvement could be more intuitively attributed to increased overall focus on informative image regions, CricaVPR’s comparatively smaller change in focus suggests that its improvements originates from other factors, possibly concerning its aggregator $g^{\phi}_{s}$.
Figure 4: Activation Maps: The selected $I^{l}$ queries (top of figure) are successfully recalled by $f_{s}$, while $f_{t}$ fails. For CricaVPR, the focus of its ICKD-trained $f_{s}$ is visualized by averaging its $z^{l}$ across the channel dimension in accordance with its authors. For NetVLAD, we averaged a weighted sum of its MSE-trained $z^{l}$ via tentative cluster assignments from $g^{\phi}_{s}$. For MixVPR, since its authors did not visualize activations, we masked small regions of $I^{l}$ and considered the resulting change in its ICKD-trained $v_{l}$ as activations, creating block-like patches on its heatmap.
Table III: Indoor Evaluation Using Pretrained Weights
Methods Quality TUM LSI Gangnam Station NYC-Indoor- VPR
R@1 R@5 R@1 R@5 R@1 R@5
MixVPR $I^{h}$ 94.09 99.55 4.39 13.63 41.23 83.76
90% 91.36 99.09 3.93 11.76 40.04 80.76
CricaVPR $I^{h}$ 92.27 99.09 7.48 21.64 37.92 82.36
90% 92.27 97.27 7.67 21.37 37.30 79.26
DINOv2 SALAD $I^{h}$ 94.09 99.09 8.93 22.67 39.89 82.98
90% 91.82 98.64 9.43 21.37 39.58 81.69
NetVLAD $I^{h}$ 95.00 99.09 3.21 8.36 39.42 81.27
90% 91.82 98.18 2.94 9.81 38.23 79.10
AnyLoc $I^{h}$ 97.73 99.55 4.47 12.79 37.92 82.77
90% 93.64 98.18 3.63 11.11 36.37 81.07
VI Additional Experiments
While our knowledge distillation approach already demonstrates notable efficacy for the majority of methods and datasets tested, we extend distillation to more datasets and other aforementioned routes of image quality reduction II-A.
VI-A JPEG Compression’s Impacts on Indoor Datasets
While our previous VPR testing datasets cover both rural and urban environments and contain temporal changes and perceptual aliasing, they all capture outdoor environments. In table III, We further examine the impacts of JPEG compression on three indoor datasets of different difficult levels, namely TUM LSI [27], Gangnam Station (one of NAVER LABS’ large-scale localization datasets [30]), and NYC-Indoor-VPR [33]. While the performance of all methods with pretrained weights fluctuate greatly across datasets, JPEG compression’s impacts on recall rates are significantly smaller than previous results with outdoor datasets. We postulate that due to indoor environments generally lack large contiguous regions of uninformative features such as sky, the proportion of distinctive features degraded by JPEG compression is smaller within indoor images. Therefore, instead of repeating our previous experiments, we extend our distillation methodology to other possible forms of $I^{l}$.
VI-B Other Modalities of Image Quality Reduction
Inferring from JPEG compression’s greater impacts on outdoor scenes, we use outdoor datasets from IV-A2 to yield $I^{l}$ with reduced resolution instead of JPEG compression. Secondly, we explore the impacts of video quantization on VPR recall, mirroring realistic practice of streaming image data through video. However, the image-based nature of existing VPR datasets prompts us to curate a custom video-based VPR dataset to produce $I^{l}$ with increased video quantization.
Amongst the five methods in table II, NetVLAD suffers the most from lowered image quality in general. Thus, we select NetVLAD as the candidate of verifying knowledge distillation’s benefits for VPR performance under resolution and quantization-based $I^{l}$.
VI-C Custom Video-Based VPR Dataset
Our custom VPR dataset is sourced from indoor spaces of the 6th floor of the Lighthouse Guild, an eye care facility in New York City. Instead of following existing indoor VPR datasets in capturing images [27, 30, 31, 33], we record panoramic videos using the Insta360 camera, providing a $360$°  view. Each frame, with dimensions $5760\times 2880$, was horizontally segmented into $18$ perspective images with $90$°  field of view and dimensions $1440\times 810$. These images were assembled into perspective videos for both resolution reduction and video quantization according to the following:
  • Resolutions: $405$p ($720\times 405$), $203$p ($360\times 203$)
  • Quantization Parameters: $30$, $33$, $36$, $39$, $42$, $45$, $48$
VI-D Resolution and Quantization Experimental Results
We not only validate distillation’s efficacy for different forms of $I^{l}$ but also its independence to datasets and $I^{l}$’s modality during training. While following IV-C1’s steps, we perform distillation on NetVLAD using the Pitts250k dataset [25], which contains around $250,000$ outdoor $640\times 480$ images of Pittsburgh’s urban environments. All images are reduced to $240\times 180$ as $I^{l}$ for $f_{s}$ to learn on. To analyze distillation’s robustness to changing $I^{l}$ modalities, we test NetVLAD’s recall on our custom dataset with $v^{l}$ extracted from resolutions of $405$p and $203$p and quantization parameters spanning from $30$ to $48$ in intervals of $3$. For previous outdoor datasets, we extract $v^{l}$ from $320\times 240$ ($240$p). The influence of our dataset’s various resolutions and quantization levels on $f_{s}$’s recall rate after distillation is compared against pretrained weights as a baseline in Fig. 5. The $203$p results restate MSE (2)’s efficacy on NetVLAD, but despite ICKD (1)’s comparatively inferior performance on constant-modality $I^{l}$ before, combining the two losses further improves VPR performance on $405$p, suggesting ICKD’s greater adaptability under video quantization. Notably, changing recall distance thresholds as defined in IV-C2 has little effect on VPR performance.
For outdoor datasets, we compare pretrained NetVLAD against trained $f_{s}$ with the highest recall rate in table IV. NetVLAD’s best-performing distillation loss remains MSE, but the addition of ICKD yields stronger performance on Tokyo 24/7. This corroborates the previous observation of ICKD’s adaptability, which may allow $f_{s}$ to better withstand changing $I^{l}$ modalities from $180$p during distillation to $240$p in testing.
Figure 5: Results on Our Dataset: The recall rates of Pitts250k-trained $f_{s}$ are obtained for each quantization parameter. For ease of visualization, the latter is converted into the video bitrate metric with lower bitrate indicating higher quantization. The left two plots were calculated with a VPR threshold of $1$m, and the right two have a $10$m threshold.
Table IV: NetVLAD’s Performance on $240$p Images
Dataset Configuration R@1 R@2 R@5 R@10
Mapillary SLS pretrained 31.16 37.89 46.36 52.54
MSE +7.57 +7.88 +7.64 +7.54
Nordland pretrained 3.99 5.40 8.34 10.40
MSE +1.12 +1.23 +1.23 +1.81
Tokyo 24/7 pretrained 24.84 29.62 39.49 48.41
ICKD + MSE +7.64 +8.28 +7.64 +7.32
VII Conclusion
Addressing the concern of reduced VPR accuracy under the realistic scenario of image quality reduction, our knowledge distillation methodology for extracting more discriminative descriptors from low quality images achieves significant performance gains under JPEG and video compression for various VPR methods and datasets. Furthermore, we have demonstrated our approach’s potential to generalize to other kinds of low quality images such as low resolutions, reducing the literature gap on video-based VPR datasets in the process.
A suitable direction of future work is to justify distillation’s effects on different kinds of VPR methods. While we have identified the general strengths of ICKD and MSE losses and noted the specific usability of the triplet loss, a better understanding of each VPR method’s affinity to specific losses could reveal further insights on VPR methods themselves. Secondly, more VPR datasets and research under the context of video quantization could prove valuable for the future real-world applications of VPR.
Acknowledgment
We would like to express our gratitude to Zezheng Li and Liyuan Geng for their valuable assistance with test data preprocessing. Their contributions helped facilitate the progress of this work.